Practice question: empirical formula for 7.615g of Ga, 2.622g of O
Answer: Ga2O3
How would we get 2 for the final answer for Ga since it would be (7.615g/1)(mol/69.7g) = 0.109 mol and would be divisible by itself since it is the lowest mole so would only equal to 1? If it were to be multiplied by something, what is it and why?
(since it's already a whole number why do we need to change it?)

- Thank you!
1
Expert's answer
2020-02-21T07:53:36-0500

n(Ga) = 7.615/69,7 = 0.109 mol

n(O) = 2.622/16 = 0.164 mol

n(Ga):n(O) = 1 : 1.5 = 2 : 3

We need to multiply this ratio by 2 because 1.5 is not integer.

Answer: Ga2O3